OpenStudy (anonymous):

What is the value of n? 12+4n / 3 = 8

OpenStudy (mathstudent55):

Is it \(\dfrac{12 + 4n}{3} = 8\) or \(12 + \dfrac{4n}{3} = 8\) ?

OpenStudy (anonymous):

The top

OpenStudy (anonymous):

same question... @mathstudent55

OpenStudy (mathstudent55):

Ok. Since the entire left side is a fraction with denominator 3, then you must multiply both sides by 3. Try that and show what you get.

OpenStudy (anonymous):

I got 8 for my answer. I dont know if thats right.

OpenStudy (mathstudent55):

\(\dfrac{12 + 4n}{3} = 8\) \(\dfrac{12 + 4n}{3} \color{red}{\times 3} = 8\color{red}{\times 3} \)

OpenStudy (mathstudent55):

I don't know if n = 8 is correct because I haven't solved the equation yet.

OpenStudy (anonymous):

I had solved it and got 8. I dont know though.

OpenStudy (mathstudent55):

The answer is definitely not 8. All you need to do is plug in 8 in the original equation, and you'll see that 8 does not work. How about trying to solve the way I'm explaining to you? The first step is to multiply both sides by 3 to get rid of the denominator. \(\dfrac{12 + 4n}{3} \color{red}{\times 3} = 8\color{red}{\times 3}\) \(12 + 4n = 24\) I did the first step for you. Both sides were multiplied by 3. Now, since there is a 12 on the left side, subtract 12 from both sides.

OpenStudy (anonymous):

Im confused haha. Im not very good at math.

OpenStudy (mathstudent55):